OpenStudy (anonymous):

converges or diverges summation symbol from 1 to infinity (0.001 + (k^(-1.5) Use divergence test

OpenStudy (turingtest):

\[\sum_{k=1}^{\infty}0.001+k^{-3/2}\]?

OpenStudy (anonymous):

\[\sum_{1}^{\infty} 0.001 + k ^{-1.5}\]

OpenStudy (anonymous):

yeah i'm not sure how to use equation editor when posting a question. i'm still new to this

OpenStudy (turingtest):

no prob, better to learn the LaTeX language when you get the chance what kind of test are you supposed to use here? there are a few "divergence tests" as far as I know

OpenStudy (turingtest):

is this the one where we take the limit of the partial sums? hm.... I'm not the best at these @satellite73 @amistre64 series problem

OpenStudy (anonymous):

i know we are suppose to take the limit of \[a_{k}\] and see if it equals zero, then we know its convergent. but thats all i know

OpenStudy (anonymous):

wouldn't your \[a_{k}\] = \[k^{-1.5}\]

OpenStudy (anonymous):

ok if, \[\lim_{k \rightarrow \infty}.001+k^{-1.5}=0 \] then it will be convergent.. so let's do that: \[\lim_{k \rightarrow \infty}.001+k^{-1.5}=\lim_{k \rightarrow \infty}(1/1000)+(1/k^{1.5})\] \[\lim_{k \rightarrow \infty} (k^{1.5}+1000)/(1000k^{1.5})\] to make things easier I'll use the L'Hopital stuff... \[=\lim_{k \rightarrow \infty} (1.5k^{.5})/(1.5)(1000k^{.5})=\lim_{k \rightarrow \infty} 1/1000=1/100\] since \[\lim_{k \rightarrow \infty} .001+k^{-1.5} \neq 0\] then the series will diverge

OpenStudy (anonymous):

@mrlunarugby the a sub n will be .001+k^(-1.5)

OpenStudy (anonymous):

so suppose if we have \[\sum_{n=1}^{\infty}(oaiwejnofivjawelravefav)^n+n^2+e^n+logn\] our a sub n will be oaiwejnofivjawelravefav)^n+n^2+e^n+logn and to find if this series diverges, we find the \[\lim_{n \rightarrow \infty} a_n=\lim_{n \rightarrow \infty} oaiwejnofivjawelravefav)^n+n^2+e^n+logn\] if \[\lim_{n \rightarrow \infty} oaiwejnofivjawelravefav)^n+n^2+e^n+logn=0\] then it is convergent. otherwise, it will be divergent

OpenStudy (anonymous):

Right. thanks it took me a minute to understand it. the way you changed 0.001 to 1/1000 then k^(-1.5) to 1/k^1.5 i just didn't think to do it that way. that was very helpful. thanks and by the way in the question that i asked you shouldn't the limit approach 1/1000 not 1/100. was that a typo??

OpenStudy (anonymous):

how do i type equations when i'm posting questions? because i have more problems i need help with.

OpenStudy (turingtest):

gotta learn the language LaTeX to type equations quickly

OpenStudy (anonymous):

ohh is that a programming language

OpenStudy (anonymous):

i'm only learning java at the moment

OpenStudy (turingtest):

it's not that hard, here's a cheat sheet: http://omega.albany.edu:8008/Symbols.html

OpenStudy (anonymous):

oh thanks. i can't learn all this know though i'm trying to cram for final

OpenStudy (anonymous):

now*

OpenStudy (turingtest):

you have to enclose things in brackets i.e. \sum_{i=1}^{\infty}i enclose this in brackets like so \[\sum_{i=1}^{\infty}i\.] ^without this dot and it becomes...

OpenStudy (turingtest):

\[\sum_{i=1}^{\infty}i\]

OpenStudy (turingtest):

but anyway, anonymous found that the series diverges because the limit of the summand as \(k\to\infty\neq0\) that is how that trick worked

OpenStudy (anonymous):

yeah that was very helpful

OpenStudy (turingtest):

mind you that if\[\lim_{k\to\infty}\{a_n\}=0\]that does \9not\) mean that the series does converge, only that it might converge

OpenStudy (turingtest):

*does not

OpenStudy (turingtest):

for example\[\sum_{i=1}^{\infty}\frac1i\]diverges\[\sum_{i=1}^{\infty}\frac1{i^2}\]converges

OpenStudy (anonymous):

@mylunarugby oh yeah :)) my bad, either it was a typo or there was not enough space :))